I don't know how to do this problem. can someone please help me..please.

"The peremeter of a rectangle is 100ft. One side is 6ft longer than the other side. An equation o find the length of "x" of the shorter side is what????"

Answer by ionesco(12)

x is the shorter side
y is longer side
peremeter is 2x + 2y =1000, meaning x + y = 500,
y longer than x y = 6x
then the first becomes x+ 6x=500, then 7x=500 then x=500/7